# Proxima: a DAG–based cooperative distributed ledger

### v4

@@ Abstract
This paper introduces a novel architecture for a distributed ledger, commonly referred to as a blockchain, which is organized in the form of a directed acyclic graph with UTXO transactions as vertices, rather than as a chain of blocks. The consensus on the state of ledger assets is achieved through cooperative consensus: a profit-driven behavior of token holders themselves, which is viable only when they cooperate by following the biggest ledger coverage rule. The cooperative behavior is facilitated by enforcing purpose-designed UTXO transaction validity constraints.
@@
The setup allows to achieve high throughput and scalability together with low transaction costs, while preserving key aspects of permissionless decentralization and asynchronicity found in Bitcoin and other proof-of-work blockchains, but without unreasonable energy consumption. Sybil protection is achieved similarly to proof-of-stake blockchains, using tokens native to the ledger, yet the architecture operates in a leaderless manner without block proposers and committee selection.333 I am grateful to Serguei Popov, Bartosz Kuśmierz and @buddhini for constructive criticism, which I believe made the whole thing better.
@@ 1 Preface
With Proxima, we make no claims about solving all blockchain problems such as trilema or being better than any other solutions. However, we hope that it will attract interest from blockchain system developers and academia. To our best knowledge at the time of writing, cooperative consensus is the only alternative to the PoW consensus principle, which preserves the trait of permissionless decentralization, a unique selling point of crypto ledgers in general. Some aspects, convergence and security in particular, certainly require deeper mathematical modeling and analysis.
@@ 2.1 Generic principles of the permissionless distributed ledger
The principle of one transaction approving two previous transactions results in a DAG-like data structure that is constructed incrementally by adding new vertices. DAG edges represent approval links. The specific number of approved transactions (two) is not essential; we generalize it to "any number from 1 to a fixed $N$".
@@ 3.1 UTXO transaction
$(in(T),out(T),endorse(T),timestamp(T),unlock(T),senderID(T))$
@@ 3.5.3 Validity of outputs
The absolute majority of transaction validity checks are expressed through the use of validation scripts, immutably attached to outputs in ledger state.
@@
The ledger definition contains a static library of standard scripts that are included in the ledger genesis. For example, the standard chain constraint is a library script, addressED25519 lock script is a standard script, and there are many others. [^18] Note, that each script is taking its parameter data from the context only. For example, running the script $addressED25519(0x97459a12301950535caa5115af5edbc04ead272c25880c967ee522dc11bc8193)$ in the consumed output means checking if the output which is locked in this particular target address is unlocked in the context of the consuming transaction.
@@ 4 Nodes
  • Protect against spamming attacks: nodes are responsible for protecting themselves and their peers from potential spamming attacks to maintain network integrity (see Spamming prevention).
@@ 6 Chains
A transaction may belong to several chains. Chain has unlimited-length [^24] The concept, known as the chain on the UTXO ledger, was first introduced in the IOTA’s Stardust ledger specification as Alias output;
Outputs in the UTXO ledger are single-time transient assets that "live" until they are consumed by a transaction. The chain transforms a sequence of outputs into a permanent, non-fungible asset with potentially unlimited lifetime. This provides an important mechanism for continuity and permanence in the ledger. Thus, the chain constraint enforces a chain of outputs on the UTXO tangle, identified by the chain ID as a unique and persistent identifier for the chain’s lifetime.
@@
Inflation is increasing the total supply of tokens in the ledger. In Proxima, inflation serves as the basis for an incentive mechanism and is closely related to the concept of chains. This design aims to reward participants in the network for maintaining the ledger’s integrity and promoting healthy economic dynamics within the ecosystem. Proxima enables the controlled generation of new tokens over time, offering a clear and predictable structure for incentives.
Let us assume that the predecessor (chained) output $O$ has timestamp $t$, and its successor output $O^{\prime}$ has timestamp $t+\Delta t$. It is obviously true that $O\leftarrow O^{\prime}$.
The creation tokens "out of thin air" is given by the formula:
$amount(O^{\prime})=amount(O)+inflation(O^{\prime})$
As a general principle, we want inflation to be proportional to time and capital:
$inflation(O^{\prime})=R\times amount(O)\times\Delta t$
where $R$ is an inflation rate.
For practical reasons, we modify the pure linear formula above with the goal of incentivizing the desired behavior of token holders.
Let us define distance in full slots between O and O’:
$d(O,O^{\prime})=\begin{cases}slot(t+\Delta t)-slot(t)&\text{if }ticks(t)\neq 0\text{ (a branch)}\\ slot(t+\Delta t)-slot(t)+1&\text{if }ticks(t)=0\text{ (not branch)}\end{cases}$
Then inflation is calculated according to the following rules:
Definition 7.1 (Inflation on the chained output).
$inflation(O^{\prime})=\begin{cases}BB&\text{ if O' is a branch}\\ d(O,O^{\prime})\times amount(O)\times R_{t}&\text{if }d(O,O^{\prime})\leq W\text{ and O' is not a branch}\\ W\times amount(O)\times R_{t}&\text{if }d(O,O^{\prime})>W\text{ and O' is not a branch}\\ \end{cases}$
Here:
- $W$ is a ledger constant called linear inflation slots252525in the testnet linear inflation slots is 3 slots ($\approx 30$ seconds).
- $R_{t}$ is an inflation rate, adjusted for the ledger time $t$ [^26] In the testnet, the inflation rate corresponds to flat inflation (tail inflation): annual  10% on the total supply the first year and the diminishing
- $BB$ is branch inflation bonus. It is a randomized value calculated for each branch transactions. (see details in Section 9.2.Rich chain bias. Meta-stability. Random branch inflation).
The enforced inflation value gives linear growth of the inflated tokens on the chain as long as the token holder issues at least one chained output per $W$ slots. If the gap between the chained outputs is longer than $W$, inflation stops growing and return-on-capital starts diminishing. Likewise, branch inflation bonus encourages token holders to issue sequencer transactions.
It is enforced that inflation on the chained output does not exceed the maximal calculated amount.
@@
The balance validity check of input-output amounts for a transaction $T$ ensures that the sum of the output amounts produced does not exceed the sum of the input amounts plus the total inflation of all produced chained outputs. This check is essential to maintain the integrity of the ledger and to enforce the inflation limits set by the chain constraints. Given a transaction $T$, the validity check can be expressed as follows:
$\sum_{o\in in(T)}amount(o)\leq\sum_{o\in out(T)}amount(o)\leq\sum_{o\in in(T)}amount(o)+\sum_{o\in out(T)}inflation(o)$
This inflation mechanism creates incentives for token holders to participate in constant chain creation and maintenance, as they can potentially gain additional tokens over time as long as they issue at least one chain-constrained output per linear inflation slots $W$.
@@
    token holders can choose to run a sequencer. By doing so, they build the sequencer chains with inflation and potentially benefit from transaction fees.
@@
    token holders earn inflation rewards through the delegation mechanism. The delegation effectively is lending tokens to support the sequencer’s operations, while expecting a share of the generated inflation.
@@ 8.1 Sequencers
There is substantial evidence that distributed ledgers tend to converge towards a high concentration of capital no matter what, where a small number of entities control the vast majority of assets on the ledger282828see for example How centralized is decentralized? Comparison of wealth distribution in coins and tokens: https://arxiv.org/abs/2207.01340. Prominent examples include major exchanges and token custodians which are entrusted to manage large amounts of on-chain assets. This appears to be a fundamental and unavoidable trend.
With Proxima, instead of worrying about "fairness" of the capital distribution or attempting to create a more equitable world, the design embraces practicalities. It assumes that a significant concentration of tokens in the hands of a few is normal and follows the power law, exhibited by many biological and social systems. Consequently, major token holders (often "wales" but primarily large exchanges and other custodians) will eventually command a substantial share of the capital, which entails a substantial amount of skin-in-the-game, commitment, and trust in the distributed ledger (more about this in 10.4.Ultimate security. Social consensus).
@@ 8.3 Sequencer baseline
    $T$ endorses (see endorsement) another transaction $T^{\prime}$ on the same slot as $T$: $T^{\prime}\Leftarrow T$. $T$ will inherit the baseline state from $T^{\prime}$ recursively.
@@
The chain of branches represents the history of ledger states since the genesis ledger state, which ends with the baseline of the transaction303030chain of branches, is similar to the chain of blocks in a blockchain.
@@ 8.4 Sequencer chain building
There are two sequencer chains: chain $A$ and chain $B$. In slot $N-1$, branch $A$ extends into two conflicting branches in slot $N$: Branch $A$ and Branch $B$. This is achieved by both chains consuming the same stem output, resulting in conflicting transactions (transactions 1 and 2) in slot N.
@@ 8.6 Ledger coverage
We will denote it by $coverage\Delta(T)$. The past cone of the sequencer transaction $T$ends" at baseline state $G_{L_{T}}$. See also Figure 3.
@@
$coverage(T)=\begin{cases}initialSupply&slot(T)=1\\ coverage\Delta(T)+\frac{coverage\Delta(baseline(T))}{2^{slot(T)-slot(baseline(T))+1}}&\text{T is not a branch}\\ coverage\Delta(T)+\frac{coverage\Delta(stemPredecessor(T))}{2^{slot(T)-slot(stemPredecessor(T))}}&\text{T is a branch}\\ \end{cases}$
@@
For practical reasons, we use integer arithmetic and assume that the values of $coverage$ fit 64-bit integers; therefore, we can restrict summing along the maximum $N=64$ slots back.
@@ 8.7 Sequencing strategy
[…]
@@
1:procedure GenericSequencer($sequencerID,privateKey,utxoTangle,localClock$)
@@
4:         $bestCandidate\leftarrow nil$triangleright assume $coverage(nil)=0$
@@
6:         while $localClock.now<timestampTarget.realTimeValue$ do
@@
9:              if $coverage(candidate)>coverage(bestCandidate)$ then
@@ 8.7.1 Pace of the sequencer chain
[…]
@@
    Transactions are subject to a minimum interval between timestamps, defined by the transaction pace validity constraint. This constraint sets a lower bound on the timestamp for the next sequencer transaction based on the given inputs. One of objectives behind pace constraints, is to put a cap on maximum transaction issuance rate per token holder (see Spamming prevention).
@@ 8.8 Non-sequencing users
Leaving tag-along output with zero or very small fees creates also a problem of dusting313131huge number of small amount UTXOs which bloat the database. We will elaborate on the solution for that problem in section 11.Spamming prevention.
@@ 9.1 Growing the UTXO tangle
Let us consider the figure 11 ($T_{i}$ will designate the transaction with the number $i$ in the picture). Let us say that sequencer milestones of the sequencers $A$ ($T_{4}$) and $B$ ($T_{9}$) at the end of the previous slot have respective token amounts $amount_{A}$ and $amount_{B}$. Furthermore, let us assume that the ledger coverages of the transactions ($T_{4}$) and ($T_{9}$) are equal. This means that the coverages of branches $A$ and $B$ are also equal:
@@
The amounts in sequencer milestones of the branches $A$ ($T_{3}$) and $B$ ($T_{8}$) are $amount_{A}+I_{A}$ and $amount_{B}+I_{B}$, where $I_{A}$ and $I_{B}$ are branch inflation bonus amounts.
@@
    where $s=VRF_{pred}||slot$ is the value of VRF of the predecessor branch transaction concatenated with the slot number, and $rnd$ is generated random value, $proof$ is the cryptographic proof of randomness. ($s$ is enforced and never repeats)
@@
  • The following constraint on values $proof$ and $rnd$ is enforced on the produced sequencer milestone:

### v5

@@ Abstract
This paper introduces a novel architecture for a distributed ledger, commonly referred to as a blockchain, which is organized in the form of a directed acyclic graph with UTXO transactions as vertices, rather than as a chain of blocks. The consensus on the state of ledger assets is achieved through cooperative consensus: an emerging profit-driven behavior of token holders themselves, which is viable only when they cooperate by following the biggest ledger coverage rule. The cooperative behavior is facilitated by enforcing purpose-designed UTXO transaction validity constraints.
@@
The setup allows to achieve high throughput and scalability together with low transaction costs, while preserving key aspects of permissionless decentralization and asynchronicity found in Bitcoin and other proof-of-work blockchains, but without huge energy consumption. Sybil protection is achieved similarly to proof-of-stake blockchains, using tokens native to the ledger, yet the architecture operates in a leaderless manner without block proposers and committee selection. [^3] I am grateful to Serguei Popov, Bartosz Kuśmierz and @buddhini for constructive criticism, which I believe made the whole thing better.
@@ 1 Preface
With Proxima, we make no claims about solving all blockchain problems such as trilema or being better than any other solutions. However, we hope that it will attract interest from blockchain system developers and academia. The cooperative consensus is one of just a few alternatives to the PoW consensus principle, which preserves the trait of permissionless decentralization, a unique selling point of crypto ledgers in general. Some aspects, convergence and security in particular, certainly require deeper mathematical modeling and analysis.
@@ 2.1 Generic principles of the permissionless distributed ledger
The principle of one transaction approving two previous transactions results in a DAG-like data structure that is constructed incrementally by adding new vertices. DAG edges represent approval links. The specific number of approved transactions (two) is not essential; we generalize it to "any number from 1 to a fixed NN".
@@ 3.1 UTXO transaction
(in(T),out(T),endorse(T),timestamp(T),unlock(T),senderID(T))(in(T),out(T),endorse(T),timestamp(T),unlock(T),senderID(T))
@@ 3.5.3 Validity of outputs
The absolute majority of transaction validity checks are expressed through the use of validation scripts, immutably attached to outputs in ledger state. The UTXO is also said to carry a covenant, a kind of (smart) contract which is enforced by protocol between the owner of the output and the consumer of it.
@@
The ledger definition contains a static library of standard scripts that are included in the ledger genesis. For example, the standard chain constraint is a library script, addressED25519 lock script is a standard script, and there are many others. [^18] Note, that each script is taking its parameter data from the context only. For example, running the script addressED25519(0x97459a12301950535caa5115af5edbc04ead272c25880c967ee522dc11bc8193)addressED25519(0x97459a12301950535caa5115af5edbc04ead272c25880c967ee522dc11bc8193) in the consumed output means checking if the output which is locked in this particular target address is unlocked in the context of the consuming transaction.
@@ 4 Nodes
  • Protect against spamming attacks: nodes are responsible for protecting themselves and their peers from potential spamming attacks to maintain network integrity (see Denial-of-service prevention).
@@ 6 Chains
A transaction may belong to several chains. The chain has unlimited length.
Outputs (UTXOs) in the UTXO ledger state are single-time transient assets that exist until they are consumed by a transaction. The chain transforms a sequence of outputs into a permanent, non-fungible asset with potentially unlimited lifetime. This provides an important mechanism for continuity and permanence in the ledger. Thus, the chain constraint enforces a chain of outputs on the UTXO tangle, identified by the chain ID as a unique and persistent identifier for the chain’s lifetime.
@@
Inflation is increasing the total supply of tokens in the ledger. In Proxima, inflation serves as the basis for an incentive mechanism and is closely related to the concept of chains. This design aims to reward the participants in the network for maintaining the integrity of the ledger and promoting healthy economic dynamics within the ecosystem with the controlled generation of new tokens over time242424current version of the inflation mechanism simplifies the one presented in the earlier versions of the whitepaper by removing thresholds and making it more linear and directly computable on-ledger with basic formulas of EasyFL.
As a general principle, we want inflation $I$ to be proportional to time and funds $A$:
$I=A\times R\times\Delta t$
where $\Delta t$ is the duration of time and $R$ is an inflation rate.
7.1 Inflation definitions
To apply this general principle to the discrete time scale and specifics of the Proxima ledger, let us consider a chain transition (produced by a transaction) where the predecessor is a (chained) output $O$ and its successor is output $O^{\prime}$.
Let us denote $A=amount(O)$ and $A^{\prime}=amount(O^{\prime})$. $I$ the number of new tokens created in this transition:
$A^{\prime}=A+I$
Let $t$ be the slot of the output $O$ and $\Delta t$ be the number of full slots between $O$ and $O^{\prime}$. We define $I_{t}$ inflation on the slot $t$ of the amount $A$ as follows:
Definition 7.1 (Single-slot inflation).
$I_{t}=\begin{cases}0&\Delta t=0\\ R_{t}\cdot A&\Delta t\geq 0\end{cases}$
Where
$R_{t}=\frac{1}{C+t}$
$R_{t}$ is the inflation rate in the slot $t$. It declines with every slot. The constant $C$ defines how steep the decline is.
The transition of a chain by a transaction generates zero new tokens if the predecessor and successor are in the same slot. The inflation amount is proportional to $R_{t}$ if it is in different slots.
Let us say $A_{t}$ is some amount in the chain output in the slot $t$. The successor of the output will always produce inflation $I_{t}$ (one slot), no matter how long the gap between the outputs is.
The sequence of $k$ chained outputs in $k$ subsequent slots will produce amount on the output:
$A_{t+k}=A_{t}\cdot(1+R_{t})(1+R_{t+1})\dots(1+R_{t+k-1})=A_{t}\cdot\prod_{i=t}^{t+k-1}(1+R_{i})$
$1+R_{i}=\frac{C+i+1}{C+i}\text{ for }i>0$
$A_{t+k}=A_{t}\cdot\frac{\bcancel{C+t+1}}{c+t}\cdot\frac{C+t+1+1}{\bcancel{C+t+1}}\dots\frac{C+t+k-1+1}{\bcancel{C+t+k-1}}=\frac{C+t+k}{C+t}$
$A_{t+k}=A_{t}\cdot(1+k\cdot R_{t})$
The formulas
$I_{A,t,k}=k\cdot\frac{A}{C+t}$
$A_{t+k}=A_{t}+k\cdot\frac{A_{t}}{C+t}$
respectively gives inflation of the amount $A$ starting with the slot $t$ over the period of $k$ slots and the resulting number of tokens.
Figure 8: Inflation rate projections C=30,303,030
This is our intention: inflation and the total amount are linear over time.
We adjusted the formulas above for integer arithmetic. We can see that for inflation $\frac{A_{t}}{C+t}$ to be positive in integer arithmetic, we need the following condition satisfied:
$A_{t}\geq C+t$
The above is the condition of minimum on-chain balance with non-zero inflation where $C$ plays the role of minimum inflatable amount of tokens in slot 0.
In figure 8, we provide projected annual inflation values with $C=30,303,030$:
@@ 7.1 Inflation definitions
The balance validity check of input-output amounts for a transaction $T$ ensures that the sum of the output amounts produced does not exceed the sum of the input amounts plus the total inflation of all produced chained outputs. This check is essential to maintain the integrity of the ledger and to enforce the inflation limits set by the constraints. Given a transaction $T$, the validity check can be expressed as follows:
$\sum_{o\in out(T)}amount(o)\leq\sum_{o\in in(T)}amount(o)+\sum_{o\in out(T)}inflation(o)$
7.2 Intended behavior
We want to incentivize constant participation of token holders with their tokens in the consensus.
The enforced inflation value gives linear growth of the inflated tokens on the chain as long as the token holder issues at least one chain output per slot.
Inflating the amount of their holdings is completely permissionless, and every token holder is entitled for the inflation rewards, while stopping issuing transaction for the token holder means implicit cost of dilution. Likewise, branch inflation bonus encourages token holders to issue branch transactions.
@@ 7.2 Intended behavior
    token holders can choose to run a sequencer. By doing so, they proactively build the sequencer chains with inflation and potentially benefit from transaction fees.
@@
    token holders earn inflation rewards through the delegation mechanism. The delegation effectively is lending tokens to support the sequencer’s operations, while expecting a share of the generated inflation. The delegation mechanism is crucial for the scalability of cooperative consensus.
@@
We want to stress that the philosophy of Proxima does not encourage passive token holding behavior, common in most blockchains, where moving tokens usually means cost of fees. We want every token to participate in the security of the ledger.
The fear of inflation, also common in crypto communities, does not have any grounds in Proxima. The chain inflation is capped, predictable, and proportional for each token holder (equity principle). The additional branch inflation bonus is available to sequencers so that they can offset their costs of running the sequencer nodes and bearing risks.
We believe that deterministic universal inflation reflects common principles of how value markets function and this can be the basis for sound tokenomic of the Proxima ledger.
@@ 8.1 Sequencers
There is substantial evidence that distributed ledgers tend to converge towards a high concentration of capital no matter what, where a small number of entities control the vast majority of assets on the ledger262626see for example How centralized is decentralized? Comparison of wealth distribution in coins and tokens: https://arxiv.org/abs/2207.01340. Prominent examples include major exchanges and token custodians which are entrusted to manage large amounts of on-chain assets. This appears to be a fundamental and unavoidable trend.
With Proxima, instead of worrying about "fairness" of the capital distribution or attempting to create a more equitable world, the design embraces practicalities. It assumes that a significant concentration of tokens in the hands of a few is normal and follows the power law, exhibited by many biological and social systems. Consequently, major token holders (often "wales" but primarily large exchanges and other custodians) will eventually command a substantial share of the capital, which entails a substantial amount of skin-in-the-game, commitment, and trust in the distributed ledger (more about this in 10.6.Ultimate security. Social agreement).
@@ 8.3 Sequencer baseline
    $T$ endorses (see endorsement) another transaction $T^{\prime}$ on the same slot as $T$: $^{\prime}\Leftarrow T. TT$ will inherit the baseline state from $T^{\prime}$ recursively.
@@
The chain of branches represents the history of ledger states since the genesis ledger state, which ends with the baseline of the transaction282828chain of branches, is similar to the chain of blocks in a blockchain.
@@ 8.4 Sequencer chain building
There are two sequencer chains: chain $A$ and chain $B$. In slot N1N-1, branch $A$ extends into two conflicting branches in slot $N$: Branch $A$ and Branch $B$. This is achieved by both chains consuming the same stem output, resulting in conflicting transactions (transactions 1 and 2) in slot N.
@@ 8.6 Ledger coverage
We will denote it by $coverage\Delta(T)$. The past cone of the sequencer transaction $T$ "ends" at baseline state $G_{L_{T}}$. See also Figure 3.
@@
[…]
@@
For practical reasons, we use integer arithmetic and assume that the values of coveragecoverage fit 64-bit integers; therefore, we can restrict summing along the maximum $N=64$ slots back.
@@ 8.7 Sequencing strategy
There are two primary tasks that define the sequencer’s strategy: $\textsc{decideTimestampTarget}(\cdot)$ and $\textsc{selectInputs}(\cdot)$ (see descriptions below).
@@
1:procedure GenericSequencer(sequencerID,privateKey,utxoTangle,localClocksequencerID,privateKey,utxoTangle,localClock)
@@
4:   $bestCandidate\leftarrow nil$ $\triangleright$ assume $coverage(nil)=0$
@@
6:   while localClock.now<timestampTarget.realTimeValuelocalClock.now<timestampTarget.realTimeValue do
@@
9:     if coverage(candidate)>coverage(bestCandidate)coverage(candidate)>coverage(bestCandidate) then
@@ 8.7.1 Pace of the sequencer chain
Function $\textsc{decideTimestampTarget}(\cdot)$ is responsible for calculating the timestamp of the next sequencer transaction, effectively setting the pace of the chain. The optimal pace, or frequency of sequencer transactions, is a matter of strategy and optimization, and the sequencer must consider the following factors.
@@
    Transactions are subject to a minimum interval between timestamps, defined by the transaction pace validity constraint. This constraint sets a lower bound on the timestamp for the next sequencer transaction based on the given inputs. One of objectives behind pace constraints, is to put a cap on maximum transaction issuance rate per token holder (see Denial-of-service prevention).
@@ 8.8 Non-sequencing users
Leaving tag-along output with zero or very small fees creates also a problem of dusting292929huge number of small amount UTXOs which bloat the database. We will elaborate on the solution for that problem in section 11.Denial-of-service prevention.
@@ 9.1 Growing the UTXO tangle
Let us consider the figure 12 ($T_{i}$ will designate the transaction with the number $i$ in the picture). Let us say that sequencer milestones of the sequencers AA ($T_{4}$) and $B$ ($T_{9}$) at the end of the previous slot have respective token amounts $amount_{A}$ and $amount_{B}$. Furthermore, let us assume that the ledger coverages of the transactions ($T_{4}$) and ($T_{9}$) are equal. This means that the coverages of branches $A$ and $B$ are also equal:
@@
The amounts in sequencer milestones of the branches AA ($T_{3}$) and $B$ ($T_{8}$) are $amount_{A}+I_{A}$ and $amount_{B}+I_{B}$, where $I_{A}$ and $I_{B}$ are branch inflation bonus amounts.
@@
    where $s=VRF_{pred}||slot$ is the value of VRF of the predecessor branch transaction concatenated with the slot number, and rndrnd is generated random value, proofproof is the cryptographic proof of randomness. ($s$ is enforced and never repeats)
@@
  • The following constraint on values proofproof and rndrnd is enforced on the produced sequencer milestone: